Please help me solve this question:
For what integral value(s) of K can the trinomail x^2 + kx +12 be factored?

Answer by Alan3354(69443)

Please help me solve this question:
For what integral value(s) of K can the trinomail x^2 + kx +12 be factored?
===============
Questions are not solved, questions are answered.
Equations are solved.
======================
k is the sum of factors of 12.
--> -1 & -12 = -13
-2 & -6 = -8
-3 & -4 = -7
=============
1 & 12 = 13
2 & 6 = 8
3 & 4 = 7